OpenStudy (swissgirl):

\[\sigma_{\text{ new }}=\sqrt{\frac{\Sigma |1.01X-1.01 \overline{X}|^2}{n-1}}\]

OpenStudy (swissgirl):

@dan815

OpenStudy (swissgirl):

\[\sigma_{\text{ new }}=1.01\sqrt{\frac{\Sigma |X- \overline{X}|^2}{n-1}}\]

OpenStudy (swissgirl):

\[\sigma_{\text{ new }}=1.01\sigma_{\text{ old }}\]

OpenStudy (dan815):

perfect lol

OpenStudy (swissgirl):

ohhhh \[\sigma_{\text{ old }}=\sqrt{\frac{\Sigma |X- \overline{X}|^2}{n-1}}\]

OpenStudy (swissgirl):

\[\sigma_{\text{ old }}=\sqrt{\frac{\Sigma |X- \overline{X}|^2}{n-1}}\] \[\sigma_{ \text{ new}}=\sqrt{\frac{\Sigma |1.01X-1.01 \overline{X}|^2}{n-1}}\] \[\sigma_{\text{ new }}=1.01\sqrt{\frac{\Sigma |X- \overline{X}|^2}{n-1}}\] \[\sigma_{\text{ new }}=1.01\sigma_{\text{ old }}\]

OpenStudy (swissgirl):

\[\sigma_{\text{ new }}= (1+(1-x_m) )\sigma_{ \text { old }}\]